OpenStudy (anonymous):

-2x+5y=5 x=4y-10 Substitution method

OpenStudy (anonymous):

-2(4y-10)+5y=5 -8y+20+5y=5 -3y=-15 y=5 x=4(5)-10 x=10

OpenStudy (anonymous):

-2(4y-10)+5y=5 -8y+20+5y=5 -3y=-15 y=5

OpenStudy (anonymous):

Thank you --- I forgot to drop down the 5y when I was solving. -.-